# Cyclic Representations of $U_{q}(\widehat{\mathfrak{sl}}_{2})$ and its Borel Subalgebras at Roots of Unity and Q-operators

### v2

@@ 1.1. Background
The key to understanding the algebraic construction and properties of the Q-operator for $U_{q}(\widehat{\mathfrak{sl}}_{2})$ (generated by $e_{i},f_{i},t_{i}^{\pm 1}$, $i\in\{0,1\}$) with $q$ generic is to consider infinite-dimensional representations of the Borel subalgebra $U_{q}(\mathfrak{b}_{+})$ generated by $e_{i},t_{i}^{\pm 1}$ [BLZ96, BLZ97]. Adopting the notation of [CVW24] , there are precisely two independent infinite-dimensional $U_{q}(\mathfrak{b}_{+})$ representations $\rho_{z}$ and $\bar{\rho}_{z}$ that are not restrictions of representations of the full algebra $U_{q}(\widehat{\mathfrak{sl}}_{2})$. Two corresponding Q-operators, ${\mathcal{Q}}(z)$ and ${\widebar{{\mathcal{Q}}}}(z)$ are defined as transfer matrices (i.e. twisted traces over the auxiliary space of products of L-operators) by choosing the auxiliary spaces as $\rho_{z}$ and $\bar{\rho}_{z}$ respectively. In this story, there are also three other representations that may be chosen as auxiliary spaces: $\nu^{\mu}_{z}$ an infinite-dimensional Verma module representation of $U_{q}(\widehat{\mathfrak{sl}}_{2})$ that depends on a complex weight $\mu$; $V_{z}$ the two dimensional evaluation representation of $U_{q}(\widehat{\mathfrak{sl}}_{2})$; and $\phi_{z}$ an infinite-dimensional representation of $U_{q}(\mathfrak{b}_{+})$ which is triangular (meaning that $e_{1}$ acts non-trivially, but $e_{0}$ acts as 0). There are two relations between these different representations that give rise to corresponding properties of transfer matrices and Q-operators. They are:
@@
(ii) A short exact sequence    00{0}${\rho_{zq}}$${\rho_{z}\otimes V_{z}}$${\rho_{zq^{-1}}}$00{0}$\scriptstyle{\iota}$$\scriptstyle{\tau}$
@@
Let us denote the transfer matrices associated with the auxiliary spaces $\nu^{\mu}_{z}$ and $V_{z}$ by $T_{\mu}(z)$ and $T(z)$ respectively. The triangularity of $\phi_{z}$ means that the corresponding L-operator is a triangular matrix, which in turn means that the transfer matrix with $\phi_{z}$ as the auxiliary space is diagonal (and in a given spin sector is proportional to the identity). Relations of the following form then follow directly from (i) and (ii) respectively (details and notation can be found in [VW20, CVW24]):
@@ 1.2. Main Results
This is the analog of property (i) above. We then go on in Theorem 3.7 to give $U_{q}(\mathfrak{b}_{+})$ short exact sequences that are analogous to (ii) above. Parallel results that give these factorization and fusion properties in terms of L-operators are then given in Propositions 3.13 and 3.14 respectively. We then go in Section 4 to use these various representations as auxiliary spaces in the definition of Q operators and transfer matrices.
@@ 1.3. Goals and Motivation
There are two key motivations behind the work. Firstly, by systemizing the approach as indicated we hope to open the way to developing the construction and properties of Q-operators for higher rank cases at $q^{N}=1$. Secondly, we intend to use these results in order to construct Q-operators for open systems at $q^{N}=1$, extending the generic-$q$ results of [CVW24] to a wider class of boundary conditions and to higher rank. As we explain in Section 5, there are reasons why it is simpler to develop such open system results in the $q^{N}=1$ case first.
@@ 2.1.2. R-Matrices $\check{R}(rr^{\prime};ss^{\prime})$
$s$$r$$r$$s$$T_{rs}\sim$$:\Omega_{rs}\rightarrow\Omega_{rs}$
@@ 3.2. Fusion
We now present three short exact sequences (SESs) for different $U_{q}(\mathfrak{b}_{+})$ representations. In analogy with the classic results of [KRS81], we refer to these properties as fusion. The SESs also involve the standard two-dimensional evaluation representation $V_{z}:\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})\rightarrow\operatorname{End}{V}$ ($z\in\mathbb{C}^{\times}$, $V=\mathbb{C}^{2}$) which we define by
(3.3) $V_{z}(e_{1})=\begin{pmatrix}0&z\\ 0&0\end{pmatrix},\quad$ $V_{z}(f_{1})=\begin{pmatrix}0&0\\ z^{-1}&0\end{pmatrix},$ $V_{z}(t_{1})=\begin{pmatrix}q&0\\ 0&q^{-1}\end{pmatrix},\quad V_{z}(z_{1})=1,$
$V_{z}(e_{0})=\begin{pmatrix}0&0\\ z&0\end{pmatrix},\quad$ $V_{z}(f_{0})=\begin{pmatrix}0&z^{-1}\\ 0&0\end{pmatrix},$ $V_{z}(t_{0})=\begin{pmatrix}q^{-1}&0\\ 0&q\end{pmatrix},\quad V_{z}(z_{0})=1.$
@@
$(i)$ 00{0}${\rho_{sq}}$${\rho_{s}\otimes V_{z_{s}}}$${\rho_{sq^{-1}}}$00{0}$\scriptstyle{\iota_{s}}$$\scriptstyle{\tau_{s}}$ $\hbox{where}\quad\iota_{s}=\begin{pmatrix}c_{s}Z\\ XZ^{-1}\end{pmatrix},\;\tau_{s}=(-\frac{q}{c_{s}}XZ^{-1},Z),$
$(ii)$ 00{0}${\bar{\rho}_{sq}}$${\bar{\rho}_{s}\otimes V_{z_{s}}}$${\bar{\rho}_{sq^{-1}}}$00{0}$\scriptstyle{\bar{\iota}_{s}}$$\scriptstyle{\bar{\tau}_{s}}$ $\hbox{where}\quad\bar{\iota}_{s}=\begin{pmatrix}\bar{c}_{s}\\ XZ^{-2}\end{pmatrix},\;\bar{\tau_{s}}=(-\frac{1}{\bar{c}_{s}}XZ^{-2},1),$
@@
$(\rho_{s}\otimes V_{z_{s}})(\Delta(x))\circ\iota_{s}=\iota_{s}\circ\rho_{sq}(x),\quad\tau_{s}\circ(\rho_{s}\otimes V_{z_{s}})(\Delta(x))=\rho_{sq^{-1}}\circ\tau_{s},\quad\forall x\in U_{q}(\mathfrak{b}_{+}).$
@@
We note that while $\Omega_{rs}$ and $V_{z}$ appearing in $(iii)$ are also full $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ representations (unlike the $\rho_{s}$ and $\bar{\rho}_{s}$ representations of (i) and (ii)), $\bar{I}_{s}$ and $\bar{T}_{s}$ are only $U_{q}(\mathfrak{b}_{+})$ intertwiners and are not $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ intertwiners.
@@ 3.3. Factorization and Fusion of L-operators
(3.4) $\check{L}_{\Pi}(z)(\Pi\otimes V_{z})\Delta(x)=(V_{z}\otimes\Pi)\Delta(x)\check{L}_{\Pi}(z)\quad\hbox{for all }x\in U_{q}(\mathfrak{b}_{+}).$
In the case of $\check{L}_{\Omega_{rs}}(z)$, the relation (3.4) also extends to all $x\in\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$. As $\Pi\otimes V_{z}$ is generically irreducible as a $U_{q}(\mathfrak{b}_{+})$ representation, this requirement fixes $\check{L}_{\Pi}(z)$ uniquely up to an arbitrary multiplicative constant. As is usual when discussing R-matrices and L-operators, it is also useful for us to consider the associated operator
@@
(3.6) ${L}_{\Pi}(z)(\Pi\otimes V_{z})\Delta(x)=(\Pi\otimes V_{z})\Delta^{op}(x){L}_{\Pi}(z),\quad\hbox{for all }x\in U_{q}(\mathfrak{b}_{+}).$
@@
$s\,$$r\,$$z$$\check{L}_{\Omega_{rs}}(z)\sim$$:\Omega_{rs}\otimes V_{z}\rightarrow V_{z}\otimes\Omega_{rs}$$z$$\check{L}_{\varphi}\sim$$:\varphi_{c}\otimes V_{z}\rightarrow V_{z}\otimes\varphi_{c}$$r\,$$z$$\check{L}_{\rho_{r}}(z)\sim$$:\rho_{r}\otimes V_{z}\rightarrow V_{z}\otimes\rho_{r}$$r\,$$z$$\check{L}_{\bar{\rho}_{r}}(z)\sim$$:\bar{\rho}_{r}\otimes V_{z}\rightarrow V_{z}\otimes\rho_{r}$
@@
(3.10) $\check{\bf L}_{\Omega_{rs}}(z)(V_{z}\otimes\Pi)\Delta(x)=(\Pi\otimes V_{z})\Delta(x)\check{\bf L}_{\Omega_{rs}}(z)\quad\hbox{for all }x\in\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2}).$
@@
$z\,$$r\,$$s$$\check{\bf L}_{\Omega_{rs}}(z)\sim$$:V_{z}\otimes\Omega_{rs}\rightarrow\Omega_{rs}\otimes V_{z}$
@@
We could prove each of these relations by using uniqueness of the associated $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ intertwiner, but instead we take a more direct linear-algebra approach that exploits the simplicity of the simple form of the L-operators given by 3.9 (again, this approach is inspired by [MLP21]).
@@
$S_{r^{\prime}s}(\chi)\{U_{r}(z),V_{r^{\prime}}(z)\}\otimes\{U_{s}(z),V_{s^{\prime}}(z)\}=\{U_{r}(z),V_{s}(z)\}\otimes\{U_{r^{\prime}}(z),V_{s^{\prime}}(z)S_{r^{\prime}s}(\chi)\},$
@@
$\check{R}(z/w):V_{z}\otimes V_{w}\rightarrow V_{w}\otimes V_{z}.$
@@ 3.3.2. Fusion of L-operators
where $\check{\bf L}_{\Omega_{ss^{\prime}}}(z)$ is the $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ intertwiner $:V_{z}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{ss^{\prime}}\otimes V_{z}$ defined above, and the coefficients are given by
@@ 4.1. Quantum Space $V^{\otimes M}$
The transfer matrix of the 6-vertex model is defined as follows:
(4.1) $T(z):=\hbox{Tr}_{V^{0}}\left(R^{0M}(z)\cdots R^{02}(z)R^{01}(z)\right):V^{\otimes M}\rightarrow V^{\otimes M}.$
Note, that in this definition we use $R=P\check{R}$ for notational simplicity. The superscripts indicate on which pair of spaces in $V^{\otimes M+1}=V^{0}\otimes V^{1}\otimes\cdots\otimes V^{M}$ the R-matrix acts. The equivalent graphical realization in terms of $\check{R}$ is simply
@@
Here, and from now on, it should be understood that there is an implied trace associated with the horizontal auxiliary space in the graphical representation that we use for transfer matrices and Q-operators.
Now we can change the auxiliary representation $V_{z}$, constructed in terms of the vector space $V$, to any of the four other representations $\rho_{r}$, $\bar{\rho}$, $\Omega_{rs}$, $\varphi_{c}$ associated with the vector space $W$. Let us start with $\rho_{r}$. Then the R-matrix $R(z/w):V_{z}\otimes V_{w}\rightarrow V_{z}\otimes V_{w}$ appearing appearing in (4.1) is replaced by the corresponding L-operator $L_{\rho_{r}}(w):\rho_{r}\otimes V_{w}\rightarrow\rho_{r}\otimes V_{w}$. In this way we define a new operator
$Q_{\rho_{r}}(w):=\hbox{Tr}_{W}\left(L_{\rho_{r}}^{M}(w)\cdots L_{\rho_{r}}^{2}(w)L_{\rho_{r}}^{1}(w)\right):V^{\otimes M}\rightarrow V^{\otimes M},$
@@
$w$$w$$w$$w$$w$$w$$s$$r$$T_{\Omega_{rs}}(w):=\hbox{Tr}_{W}\left(L_{\Omega_{rs}}^{M}(w)\cdots L_{\Omega_{rs}}^{2}(w)L_{\Omega_{rs}}^{1}(w)\right)$
@@
$T_{\varphi}=q^{-(S_{z}+M)/2}\hbox{Tr}_{W}(Z^{-S_{z}}),\quad\hbox{where}\quad S_{z}=\sum_{i=1}^{M}\sigma_{z}^{i},\quad\hbox{with}\quad\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
We now present various relations between these T and Q operators that follow directly from the properties of L-operators obtained in Section 3. First of all, we note that the operators $Q_{\rho_{r}}(w)$, $Q_{\bar{\rho}_{r}}(w)$ and $T_{\Omega_{rs}}(w)$ do not commute with the total spin operator $S_{z}$: it is a consequence of the relations $X^{N}=\mathbb{I}$ that these operators only conserve spin modulo $N$. This feature of cyclic representations was noted in [BS90].
@@
The statement follows from Proposition 3.11 applied to the two auxiliary spaces in each case. ∎
@@
This follows directly from Proposition 3.13. The graphical proof is simply
@@
Relations $(i)$ and $(ii)$ both contain the same factors $C_{1}^{M}(s,w)=(b(z_{s}/w)/q)^{M}$ and $C_{2}^{M}(s,w)=(qa(z_{s}/w))^{M}$. These are precisely the coefficients those that occur in the standard TQ relations of the 6-vertex model - which are in turn used to obtain the Bethe equations (see for example [JMS21]). The apparent difference with the standard TQ relations is the more complicated dependence of the Q operators in (i) and (ii) on $s\in{\mathcal{C}}_{k}$ and $w$. However, the TQ relation may be brought into the standard form by making use of the following ‘gauge transformation’ of the matrices $U_{s}(w)$ and $V_{s}(w)$ defined in Equation (3.8):
$U_{s}(w)$ $=$ $w\,\aleph_{s}U(z_{s}/w,\mu_{s})\aleph_{s}^{-1},\quad\hbox{where}\quad U(z,\mu):=\begin{pmatrix}1&z\\ z&1\end{pmatrix}\begin{pmatrix}1&0\\ 0&\mu\end{pmatrix}\quad\hbox{and}\quad\aleph_{s}:=\begin{pmatrix}1&0\\ 0&\sqrt{\frac{\kappa_{1}y_{s}}{\kappa_{0}x_{s}}}\end{pmatrix},\quad$
$V_{s}(w)$ $=$ $w\,\beth_{s}\,V(z_{s}/w,\mu_{s})\beth_{s}^{-1}\quad\hbox{where}\quad V(z,\mu):=\begin{pmatrix}1&0\\ 0&\mu\end{pmatrix}\begin{pmatrix}-q&z\\ z&-1\end{pmatrix}\quad\hbox{and}\quad\beth_{s}:=\begin{pmatrix}1&0\\ 0&\sqrt{\frac{\kappa_{1}x_{s}}{\kappa_{0}y_{s}}}\end{pmatrix}.$

### v3

@@ 1.1. Background
The key to understanding the algebraic construction and properties of the Q-operator for $U_{q}(\widehat{\mathfrak{sl}}_{2})$ (generated by $e_{i},f_{i},t_{i}^{\pm 1}$, $i\in\{0,1\}$) with $q$ generic is to consider infinite-dimensional representations of the Borel subalgebra $U_{q}(\mathfrak{b}_{+})$ generated by $e_{i},t_{i}^{\pm 1}$ [BLZ96, BLZ97]. Adopting the notation of [CVW24] , there are precisely two independent infinite-dimensional $U_{q}(\mathfrak{b}_{+})$ representations $\rho_{z}$ and $\bar{\rho}_{z}$ that are not restrictions of representations of the full algebra $U_{q}(\widehat{\mathfrak{sl}}_{2})$. Two corresponding Q-operators, ${\mathcal{Q}}(z)$ and ${\widebar{{\mathcal{Q}}}}(z)$ are defined as transfer matrices (i.e. twisted traces over the auxiliary space of products of L-operators) by choosing the auxiliary spaces as $\rho_{z}$ and $\bar{\rho}_{z}$ respectively. In this story, there are also three other representations that may be chosen as auxiliary spaces: $\nu^{\mu}_{z}$ an infinite-dimensional Verma module representation of $U_{q}(\widehat{\mathfrak{sl}}_{2})$ that depends on a complex weight $\mu$; $\pi_{z}$ the two dimensional evaluation representation of $U_{q}(\widehat{\mathfrak{sl}}_{2})$; and $\phi_{z}$ an infinite-dimensional representation of $U_{q}(\mathfrak{b}_{+})$ which is triangular (meaning that $e_{1}$ acts non-trivially, but $e_{0}$ acts as 0). There are two relations between these different representations that give rise to corresponding properties of transfer matrices and Q-operators. They are:
@@
(ii) A short exact sequence    00{0}${\rho_{zq}}$${\rho_{z}\otimes\pi_{z}}$${\rho_{zq^{-1}}}$00{0}$\scriptstyle{\iota}$$\scriptstyle{\tau}$
@@
Let us denote the transfer matrices associated with the auxiliary spaces $\nu^{\mu}_{z}$ and $\pi_{z}$ by $T_{\mu}(z)$ and $T(z)$ respectively. The triangularity of $\phi_{z}$ means that the corresponding L-operator is a triangular matrix, which in turn means that the transfer matrix with $\phi_{z}$ as the auxiliary space is diagonal (and in a given spin sector is proportional to the identity). Relations of the following form then follow directly from (i) and (ii) respectively (details and notation can be found in [VW20, CVW24]):
@@ 1.2. Main Results
This is the analog of property (i) above. We then go on in Theorem 3.7 to give $U_{q}(\mathfrak{b}_{+})$ short exact sequences that are analogous to (ii) above. Parallel results that give these factorization and fusion properties in terms of L-operators are then given in Propositions 3.13 and 3.14 respectively. In Section 4 we use these various representations as auxiliary spaces in the definition of Q operators and transfer matrices.
@@ 1.3. Goals and Motivation
There are two key motivations behind the work. Firstly, by systemizing the approach as indicated, we hope to open the way to developing the construction and properties of Q-operators for higher rank cases at $q^{N}=1$. Secondly, we intend to use these results in order to construct Q-operators for open systems at $q^{N}=1$, extending the generic-$q$ results of [CVW24] to a wider class of boundary conditions and to higher rank. The finite-dimensionality of the $U_{q}(\mathfrak{b}_{+})$ representations involved at $q^{N}=1$ will hopefully make this a more tractable problem that in generic $q$ case.
@@ 2.1.2. R-Matrices $\check{R}(rr^{\prime};ss^{\prime})$
$s$$r$$r$$s$$T_{rs}\sim$$:\Omega_{rs}\rightarrow\Omega_{sr}$
@@ 3.2. Fusion
We now present three short exact sequences (SESs) for different $U_{q}(\mathfrak{b}_{+})$ representations. In analogy with the classic results of [KRS81], we refer to these properties as fusion. The SESs also involve the standard two-dimensional evaluation representation $\pi_{z}:\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})\rightarrow\operatorname{End}{V}$ ($z\in\mathbb{C}^{\times}$, $V=\mathbb{C}^{2}$) which we define by
(3.3) $\pi_{z}(e_{1})=\begin{pmatrix}0&z\\ 0&0\end{pmatrix},\quad$ $\pi_{z}(f_{1})=\begin{pmatrix}0&0\\ z^{-1}&0\end{pmatrix},$ $\pi_{z}(t_{1})=\begin{pmatrix}q&0\\ 0&q^{-1}\end{pmatrix},\quad\pi_{z}(z_{1})=1,$
$\pi_{z}(e_{0})=\begin{pmatrix}0&0\\ z&0\end{pmatrix},\quad$ $\pi_{z}(f_{0})=\begin{pmatrix}0&z^{-1}\\ 0&0\end{pmatrix},$ $\pi_{z}(t_{0})=\begin{pmatrix}q^{-1}&0\\ 0&q\end{pmatrix},\quad\pi_{z}(z_{0})=1.$
@@
$(i)$ 00{0}${\rho_{sq}}$${\rho_{s}\otimes\pi_{z_{s}}}$${\rho_{sq^{-1}}}$00{0}$\scriptstyle{\iota_{s}}$$\scriptstyle{\tau_{s}}$ $\hbox{where}\quad\iota_{s}=\begin{pmatrix}c_{s}Z\\ XZ^{-1}\end{pmatrix},\;\tau_{s}=(-\frac{q}{c_{s}}XZ^{-1},Z),$
$(ii)$ 00{0}${\bar{\rho}_{sq}}$${\bar{\rho}_{s}\otimes\pi_{z_{s}}}$${\bar{\rho}_{sq^{-1}}}$00{0}$\scriptstyle{\bar{\iota}_{s}}$$\scriptstyle{\bar{\tau}_{s}}$ $\hbox{where}\quad\bar{\iota}_{s}=\begin{pmatrix}\bar{c}_{s}\\ XZ^{-2}\end{pmatrix},\;\bar{\tau_{s}}=(-\frac{1}{\bar{c}_{s}}XZ^{-2},1),$
@@
$\quad(\rho_{s}\otimes\pi_{z_{s}})(\Delta(x))\circ\iota_{s}=\iota_{s}\circ\rho_{sq}(x),\quad\tau_{s}\circ(\rho_{s}\otimes\pi_{z_{s}})(\Delta(x))=\rho_{sq^{-1}}\circ\tau_{s},\quad\forall x\in U_{q}(\mathfrak{b}_{+}).$
@@
We note that while $\Omega_{rs}$ and $\pi_{z}$ appearing in $(iii)$ are also full $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ representations (unlike the $\rho_{s}$ and $\bar{\rho}_{s}$ representations of $(i)$ and $(ii)$), $\bar{I}_{s}$ and $\bar{T}_{s}$ are only $U_{q}(\mathfrak{b}_{+})$ intertwiners and are not $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ intertwiners.
@@ 3.3. Factorization and Fusion of L-operators
(3.4) $\check{L}_{\Pi}(z)(\Pi\otimes\pi_{z})\Delta(x)=(\pi_{z}\otimes\Pi)\Delta(x)\check{L}_{\Pi}(z)\quad\hbox{for all }x\in U_{q}(\mathfrak{b}_{+}).$
In the case of $\check{L}_{\Omega_{rs}}(z)$, the relation (3.4) also extends to all $x\in\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$. As $\Pi\otimes\pi_{z}$ is generically irreducible as a $U_{q}(\mathfrak{b}_{+})$ representation, this requirement fixes $\check{L}_{\Pi}(z)$ uniquely up to an arbitrary multiplicative constant. As is usual when discussing R-matrices and L-operators, it is also useful for us to consider the associated operator
@@
(3.6) ${L}_{\Pi}(z)(\Pi\otimes\pi_{z})\Delta(x)=(\Pi\otimes\pi_{z})\Delta^{op}(x){L}_{\Pi}(z),\quad\hbox{for all }x\in U_{q}(\mathfrak{b}_{+}).$
@@
$s\,$$r\,$$z$$\check{L}_{\Omega_{rs}}(z)\sim$$:\Omega_{rs}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\Omega_{rs}$$z$$\check{L}_{\varphi}\sim$$:\varphi_{c}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\varphi_{c}$$r\,$$z$$\check{L}_{\rho_{r}}(z)\sim$$:\rho_{r}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\rho_{r}$$r\,$$z$$\check{L}_{\bar{\rho}_{r}}(z)\sim$$:\bar{\rho}_{r}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\rho_{r}$
@@
(3.10) $\check{\bf L}_{\Omega_{rs}}(z)(\pi_{z}\otimes\Pi)\Delta(x)=(\Pi\otimes\pi_{z})\Delta(x)\check{\bf L}_{\Omega_{rs}}(z)\quad\hbox{for all }x\in\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2}).$
@@
$z\,$$r\,$$s$$\check{\bf L}_{\Omega_{rs}}(z)\sim$$:\pi_{z}\otimes\Omega_{rs}\rightarrow\Omega_{rs}\otimes\pi_{z}$
@@
We could prove each of these relations by using uniqueness of the associated $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ intertwiner, but instead we take a more direct linear-algebra approach that exploits the simplicity of the form of the L-operators given by 3.9 (again, this approach is inspired by [MLP21]).
@@
$S_{r^{\prime}s}(\chi)\{U_{r}(z),V_{r^{\prime}}(z)\}\otimes\{U_{s}(z),V_{s^{\prime}}(z)\}=\{U_{r}(z),V_{s}(z)\}\otimes\{U_{r^{\prime}}(z),V_{s^{\prime}}(z)\}S_{r^{\prime}s}(\chi),$
@@
$\check{R}(z/w):\pi_{z}\otimes\pi_{w}\rightarrow\pi_{w}\otimes\pi_{z}.$
@@ 3.3.2. Fusion of L-operators
where $\check{\bf L}_{\Omega_{ss^{\prime}}}(z)$ is the $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ intertwiner $:\pi_{z}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{ss^{\prime}}\otimes\pi_{z}$ defined above, and the coefficients are given by
@@ 4.1. Quantum Space $V^{\otimes M}$
The twisted transfer matrix of the 6-vertex model is defined as follows:
(4.1) $T(z):=\hbox{Tr}_{V^{0}}\left((\pi_{z}(t_{1}^{\alpha})\otimes\mathbb{I})R^{0M}(z)\cdots R^{02}(z)R^{01}(z)\right):V^{\otimes M}\rightarrow V^{\otimes M}.$
Note, that in this definition we use $R=P\check{R}$ for notational simplicity. The superscripts indicate on which pair of spaces in $V^{\otimes M+1}=V^{0}\otimes V^{1}\otimes\cdots\otimes V^{M}$ the R-matrix acts. We will discuss the twist parameter $\alpha$ below. The graphical realization in terms of $\check{R}$ that we use is simply
@@
Here, and from now on, it should be understood that there is an implied trace associated with the horizontal auxiliary space in the graphical representation that we use for transfer matrices and Q-operators. For clarity, also suppress the twist operator $\pi_{z}(t_{1}^{\alpha})$ that acts on the horizontal auxiliary space from the pictures.
Now we can change the auxiliary representation $\pi_{z}$, constructed in terms of the vector space $V$, to any of the four other representations $\rho_{r}$, $\bar{\rho}$, $\Omega_{rs}$, $\varphi_{c}$ associated with the vector space $W$. Let us start with $\rho_{r}$. Then the R-matrix $R(z/w):\pi_{z}\otimes\pi_{w}\rightarrow\pi_{z}\otimes\pi_{w}$ appearing appearing in (4.1) is replaced by the corresponding L-operator $L_{\rho_{r}}(w):\rho_{r}\otimes\pi_{w}\rightarrow\rho_{r}\otimes\pi_{w}$. In this way we define a new operator
$Q_{\rho_{r}}(w):=\hbox{Tr}_{W}\left((\rho_{r}(t_{1}^{\alpha})\otimes\mathbb{I})L_{\rho_{r}}^{M}(w)\cdots L_{\rho_{r}}^{2}(w)L_{\rho_{r}}^{1}(w)\right):V^{\otimes M}\rightarrow V^{\otimes M},$
@@
Let us now discuss the reason for including the twist parameter $\alpha$ in this and other definitions. First, we recall that in the case of generic $q$, we are forced to introduced such a twist in order to regularise the infinite-dimensional trace involved in the definition of the Q-operator (see for example, [VW20]). The cyclic representations in this paper are of course $N$-dimensional, and at first sight it would appear that the twist is unnecessary. However, with cyclic representations we encounter a different issue: if we choose the representation of the matrices $X,Z\in GL(N,\mathbb{C})$ to be $X_{ij}=\delta_{i,\hbox{mod}(j+1,N)}$, $Z_{ij}=\delta_{i,j}q^{i}$, for $i,j\in\{0,1,\cdots,N-1\}$, then it follows from basic properties of primitive roots of unity that we have
$\hbox{Tr}_{W}\left(X^{n}Z^{m}\right)=N\delta_{n\equiv 0}\delta_{m\equiv 0},$
where $\delta_{n\equiv 0}$ is defined to be equal to 1 if $n=0$ mod$(N)$, and 00 otherwise. Requiring $n\equiv 0$ is equivalent to ‘charge conservation mod $N$’ which is to be expected for transfer matrices associated with cyclic representations. Requiring that $m=0$ however means that the corresponding trace vanishes on all spin sectors apart from 00 mod $N$. The twisted trace is introduced in order to stop this phenomenon. Using the definition $\rho_{r}(t_{1})=q\mu_{r}Z^{2}$, we see that, provided we choose $2\alpha\notin\mathbb{Z}$, we have
$\hbox{Tr}_{W}\left(\rho_{r}(t_{1}^{\alpha})X^{n}Z^{m}\right)=\delta_{n\equiv 0}\;q\mu_{r}\sum\limits_{p=0}^{N-1}q^{(2\alpha+m)p}=\delta_{n\equiv 0}\;q\mu_{r}\frac{1-q^{2\alpha N}}{1-q^{2\alpha+m}}.$
[…]
@@
$w$$w$$w$$w$$w$$w$$s$$r$$T_{\Omega_{rs}}(w):=\hbox{Tr}_{W}\left((\Omega_{rs}(t_{1}^{\alpha})\otimes\mathbb{I})L_{\Omega_{rs}}^{M}(w)\cdots L_{\Omega_{rs}}^{2}(w)L_{\Omega_{rs}}^{1}(w)\right)$
@@
$T_{\varphi_{c}}=q^{-(S_{z}+M)/2}\left(\frac{q}{c}\right)^{\alpha}\frac{1-q^{2\alpha N}}{1-q^{2\alpha-S_{z}}},$
$\hbox{where}\quad S_{z}=\sum_{i=1}^{M}\sigma_{z}^{i},\quad\hbox{with}\quad\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
We now present various relations between these T and Q operators that follow directly from the properties of L-operators obtained in Section 3. First of all, we note that the operators $Q_{\rho_{r}}(w)$, $Q_{\bar{\rho}_{r}}(w)$ and $T_{\Omega_{rs}}(w)$ do not commute with the total spin operator $S_{z}$; it is a consequence of the relations $X^{N}=\mathbb{I}$ that these operators only conserve spin modulo $N$ as we have explained above. This feature of cyclic representations was noted in [BS90].
@@
The statement follows from Proposition 3.11 applied to the two auxiliary spaces in each case, along with the properties
$\check{L}_{\rho_{r}}(z)\circ\left(\rho_{r}\otimes\pi_{z}\right)\Delta(t_{1})=\left(\pi_{z}\otimes\rho_{r}\right)\Delta(t_{1})\circ\check{L}_{\rho_{r}}(z),$
$\check{L}_{\bar{\rho}_{r}}(z)\circ\left(\rho_{r}\otimes\pi_{z}\right)\Delta(t_{1})=\left(\pi_{z}\otimes\bar{\rho}_{r}\right)\Delta(t_{1})\circ\check{L}_{\bar{\rho}_{r}}(z)$
for the twist factors, which follow from the fact that the L-operators are $U_{q}(\mathfrak{b}_{+})$ intertwiners. ∎
@@
This follows directly from Proposition 3.13 and Theorem 3.3. The graphical proof (with suppression of twist insertions) is simply
@@
Relations $(i)$ and $(ii)$ of Proposition 4.3 both contain the same factors $C_{1}^{M}(s,w)=(b(z_{s}/w)/q)^{M}$ and $C_{2}^{M}(s,w)=(qa(z_{s}/w))^{M}$. These are precisely the coefficients those that occur in the standard TQ relations of the 6-vertex model - which are in turn used to obtain the Bethe equations (see for example [JMS21]). The apparent difference with the standard TQ relations is the more complicated dependence of the Q operators in $(i)$ and $(ii)$ on $s\in{\mathcal{C}}_{k}$ and $w$. However, the TQ relation may be brought into the standard form by making use of the following ‘gauge transformation’ of the matrices $U_{s}(w)$ and $V_{s}(w)$ defined in Equation (3.8):
$U_{s}(w)$ $=$ $w\,\aleph_{s}U(z_{s}/w,\mu_{s})\aleph_{s}^{-1},\;\;\hbox{where}\;\;U(z,\mu):=\begin{pmatrix}1&z\\ z&1\end{pmatrix}\begin{pmatrix}1&0\\ 0&\mu\end{pmatrix}\;\;\hbox{and}\;\;\aleph_{s}:=\begin{pmatrix}1&0\\ 0&\sqrt{\frac{\kappa_{1}y_{s}}{\kappa_{0}x_{s}}}\end{pmatrix},$
$V_{s}(w)$ $=$ $w\,\beth_{s}\,V(z_{s}/w,\mu_{s})\beth_{s}^{-1},\;\;\hbox{where}\;\;V(z,\mu):=\begin{pmatrix}1&0\\ 0&\mu\end{pmatrix}\begin{pmatrix}-q&z\\ z&-1\end{pmatrix}\;\;\hbox{and}\;\;\beth_{s}:=\begin{pmatrix}1&0\\ 0&\sqrt{\frac{\kappa_{1}x_{s}}{\kappa_{0}y_{s}}}\end{pmatrix}.$